OpenStudy (tropical):

(2 • 0 • 10^-4) • (3.1 • 10^-20)?

OpenStudy (tkhunny):

Is that multiplication by 0, in there?

OpenStudy (tropical):

The • is multiply

OpenStudy (tropical):

Sorry its 2.0 not 2 • 0

OpenStudy (tkhunny):

Aw! It was so much easier the other way. Just multiply the values and add the exponents on 10. Example: 2*10^3 * 3*10^4 = (2*3) * 10^(3+4) = 6 * 10^7